How would you prepare a 25% by mass water in ethyl alcohol using 100 ml of ethyl alcohol?
1
Expert's answer
2018-08-28T10:07:08-0400
Find the mass of 100 ml of ethyl alcohol (density = 0.789 g/ml) :
m = 100 × 0.789 = 78.9 (g)
The task is to prepare 25% by mass solution of water in ethyl alcohol, so the mass of the alcohol is 75% of the total mass. Find the mass of water required (25%) :
mw = (78.9×0.25)/0.75 = 26.3 (g)
Answer
To prepare a 25% by mass water in ethyl alcohol using 100 ml of ethyl alcohol you would add 26.3 g of water into alcohol.